# thanks: Equal contributionthanks: Equal contribution

### v1

@@ Appendix F Comparison between selection and drift dynamics
such that $R=1$ is the crossover boundary between selection and drift dominated regimes.
@@
Since we already know that our results reduce back to those in population genetics when $\alpha\ll 1$, we now focus on the case of large $\eta$ and $\alpha\gg 1$. In this regime, the solutions to $R=1$ have different types of behavior depending on the value of $s$. Recall that we require $0\leq f\leq 1$. When $s<\sqrt{4\eta/N_{\mathrm{eff}}}$ (or $\tilde{s}<2/\alpha$), there is only one solution with $s-\eta f<0$:
$f_{R=1}\sim\frac{2}{N_{\mathrm{eff}}|s|}\left(\sqrt{1+\frac{4\eta}{N_{\mathrm{eff}}s^{2}}}-\mathrm{sgn}(s)\right)^{-1}\,.$ (138)
When $s>\eta$, there is also only one solution but with $s-\eta f>0$:
$f_{R=1}\sim\frac{2}{N_{\mathrm{eff}}s}\left(\sqrt{1-\frac{4\eta}{N_{\mathrm{eff}}s^{2}}}+1\right)^{-1}\,.$ (139)
@@
Therefore, the dynamics of $f$ is qualitatively the same as the one in population genetics for $s<\sqrt{4\eta/N_{\mathrm{eff}}}$ or $s>\eta$.
@@
In contrast, the dynamics of $f$ is qualitatively distinct for $\sqrt{4\eta/N_{\mathrm{eff}}}<s\lesssim\eta$. We see that there are three solutions to $R=1$:
$f_{1}$ $=\frac{2}{N_{\mathrm{eff}}s}\left(1+\sqrt{1-\frac{4\eta}{N_{\mathrm{eff}}s^{2}}}\right)^{-1},$ (140)
$f_{2}$ $=\frac{2}{N_{\mathrm{eff}}s}\left(1-\sqrt{1-\frac{4\eta}{N_{\mathrm{eff}}s^{2}}}\right)^{-1},$ (141)
$f_{3}$ $=\frac{2}{N_{\mathrm{eff}}s}\left(\sqrt{1+\frac{4\eta}{N_{\mathrm{eff}}s^{2}}}-1\right)^{-1}\,$, (142)
with $f_{1}<f_{2}<f_{3}$. Note that we have $f_{2}\simeq f_{3}\simeq s/\eta$.
@@
As demonstrated in the main text, our results in fixation probability and mean absorption/extinction time can be qualitatively understood using the nonlinear behavior of $R$. Suppose a mutant starts with $f=f_{0}\ll 1$ at the beginning. It must overcome the first barrier $f=f_{1}$ or it becomes extinct. After that, selection brings the mutant to the region $f_{2}<f<f_{3}$, which is the coexistence region. Note that the deterministic fixed point $f^{*}\simeq s/\eta$ is within this region. Since $\left<df/dt\right>>0$ when $f<f_{2}$ and $\left<df/dt\right><0$ when $f>f_{3}$, both boundaries of the coexistence region are reflecting and the mutant is trapped inside this region. It can escape this region and reach near the absorbing boundaries $f=0,1$ only when the demographic noise reaches the tail of its distribution with exponentially low probability.
@@
Now we see that if $s<s_{c}\simeq\eta/2$, the coexistence region is closer to $f=0$ and it is exponentially more likely for the demographic noise to drive the mutant to extinction. Therefore, $p_{\mathrm{fix}}$ is exponentially suppressed for $s<s_{c}$. In contrast, when $s>s_{c}$, the coexistence region is closer to $f=1$ and it is exponentially more likely for the demographic noise to drive the mutant to fixation. The mutant becomes extinct mostly because it cannot overcome the first barrier $f=f_{1}$, hence $p_{\mathrm{fix}}$ agrees with Kimura’s formula for $s>s_{c}$.
@@
The above argument also explains the behavior of mean absorption/extinction time. It grows and decays exponentially in the region $\sqrt{4\eta/N_{\mathrm{eff}}}<s\lesssim\eta$ due to the required time to escape the coexistence region. More precisely, the time grows exponentially as $s$ increases and the coexistence region becomes farther from $f=0$. The growth continues till $s=s_{c}$. For $s>s_{c}$, the coexistence region becomes closer to fixation, and the mean absorption time decays exponentially as fixation becomes more likely. For the mean extinction time, although the time for escaping from the coexistence region to $f=0$ is even longer, it also becomes exponentially unlikely that the mutant becomes extinct after reaching the coexistence region. Namely, a mutant becomes extinct mainly due to the first barrier at $f=f_{1}$. As a result, the mean extinction time reaches maximum at $s=s_{c}$ and decays exponentially for $s>s_{c}$.
Appendix G Mean extinction time
@@
For the regimes without the coexistence region, i.e. $s<\sqrt{4\eta/N_{\mathrm{eff}}}$ or $s>\eta$, the extinction time can be approximated by the time required to reach the barrier at $f=f_{R=1}$; see Eqs. (138) and (139). For $s<\sqrt{4\eta/N_{\mathrm{eff}}}$, the barrier is reflecting and the mutant becomes extinct in the same order of magnitude of time after reaching the barrier. For $s>\eta$, the barrier is absorbing and the mutant can no longer become extinct after passing through the barrier. Now, the time for reaching the barrier can be estimated using Eq. (134); since $\Delta t$ increases with $f$, the required time to reach $f=f_{R=1}$ is dominated by $\Delta t(f_{R=1})$. Therefore, for $s<\sqrt{4\eta/N_{\mathrm{eff}}}$, the mean extinction time $T$ is
$T\sim\frac{2}{|s|}\left(\sqrt{1+\frac{4\eta}{N_{\mathrm{eff}}s^{2}}}-\mathrm{sgn}(s)\right)^{-1}\,$, (143)
and for $s>\eta$,
$T\sim\frac{2}{s}\left(\sqrt{1-\frac{4\eta}{N_{\mathrm{eff}}s^{2}}}+1\right)^{-1}\,.$ (144)
Both cases match the expectation from population genetics that $T\sim 1/|s|$ asymptotically.
On the other hand, the mean extinction time deviates from the expectation from population genetics when the coexistence region is present, i.e. when $\sqrt{4\eta/N_{\mathrm{eff}}}<s<\eta$. In this regime, the mean extinction time is dominated by the exponentially long time to escape the coexistence region driven by the stochastic dynamics. The scale of the escape time can be estimated by translating the stochastic dynamics into 1D diffusion in the presence of a potential $V(f)$. Since $f$ relaxes to near equilibrium around the deterministic fixed point $f^{*}\simeq\tilde{s}$ in the long term (but before being absorbed at $f=0,1$), the probability distribution $p(f)$ can be obtained from the Fokker-Planck equation:
$\frac{\partial p}{\partial t}$ $=-\frac{\partial}{\partial f}\left(f(1-f)(s-\eta f)p\right)+\frac{\partial^{2}}{\partial^{2}f}\left(\frac{f(1-f)}{2N_{\mathrm{eff}}}p\right)$ (145)
@@
$T_{0}\sim\exp\left(\frac{V(0)-V(\tilde{s})}{k_{B}T_{\mathrm{eff}}}\right)=\exp\left(\alpha^{2}\tilde{s}^{2}\right)\,.$ (150)
@@
$p_{0}\sim\frac{p(0)}{p(0)+p(1)}\sim\begin{cases}1&s<s_{c}\\ \frac{p(0)}{p(1)}=\exp\left(\alpha^{2}(1-\tilde{s})^{2}-\alpha^{2}\tilde{s}^{2}\right)&s>s_{c}\,.\end{cases}$ (151)
@@
$T\sim p_{0}T_{0}\sim\begin{cases}\exp\left(\alpha^{2}\tilde{s}^{2}\right)&s<s_{c}\\ \exp\left(\alpha^{2}(1-\tilde{s})^{2}\right)&s>s_{c}\,.\end{cases}$ (152)
In particular, the mean extinction time reaches maximum when $s=s_{c}\simeq\eta/2$, i.e.
@@
Appendix H Methods, including details of simulations
@@
After that, we randomly choose one of the surviving species with equal probability as the parent. Each time, we introduce a mutant associated with the chosen parent, sampled as described in Sec. A and B. We then solve the full differential equations with demographic noise using a generalized Euler method as explained below. Due to the multiplicative nature of the noise, the simulation can become numerically unstable if we sample the noise directly. Instead, we make use of the result in [68]. Within a time step $dt$, we can approximate that the total growth rates of the species remain unchanged, since the growth rates involve averages over all the species or resource abundances. In other words, the species dynamics within time $dt$ becomes
@@
To obtain fixation probabilities, we sample $10$ mutants with the same chosen parent. For each mutant, we adjust the value of $r_{m}$ (for generalized Lotka-Volterra model) or $m_{m}$ (for consumer-resource models) so that the invasion fitness difference is at a given value of $s$. We then run the above simulation $10^{3}$ times with different instances of demographic noise and count the number of fixations. We take the average of the above results to find the fixation probability $p_{\mathrm{fix}}(s)$.
@@
Appendix I Parameters used in the figures
Throughout all figures, the initial mutant frequency is $f_{0}=0.09$.
@@ Appendix I Parameters used in the figures
$S=50,\mu=2.0,\sigma=0.4,\mu_{r}=6.0,\sigma_{r}=0.6,\gamma=1,\rho=0.99,s=0.2,N_{p}=5.0,D=0.2\,.$ (157)
@@
$\sigma=0.4,\rho=0.9,N_{p}=3.3$ (159)
We change the control parameter $\alpha$ by using $D=1.0$ in Fig. 2(a) and $D=0.05$ in Fig. 2(b). In Fig. 2(e), we use
$\sigma=0.4,s=0,N_{p}=4.8\,.$ (160)
In Fig. 2(f), the simulation parameters for the case of low packing are the same as those for Fig. 2(b). To control the extent of packing, we keep all the parameters unchanged except using $\sigma=0.6$ for the case with high packing.

### v2

@@ G.2 With ecological effects
such that drift (selection) dominates the dynamics when $R<1$ ($R>1$), and $R=1$ is the crossover boundary between selection and drift dominated regimes. This is the ratio as shown in Fig. 3f-h in the main text.
Figure S5: Phase portraits of frequency dynamics in terms of selection-drift ratio. The selection-drift ratios $R$ (Eq. (140)) are calculated in different regimes of invasion fitness $s_{\mathrm{inv}}$, namely (a) $s_{\mathrm{inv}}<\sqrt{4\eta/N_{\mathrm{eff}}}$, (b) $\sqrt{4\eta/N_{\mathrm{eff}}}<s_{\mathrm{inv}}<\eta$, and (c) $s_{\mathrm{inv}}>\eta$. When $s_{\mathrm{inv}}$ is in the intermediate range, there are more crossovers between selection and drift dominated regimes, hence a coexistence region centered at $f^{*}$ emerges. The selection-drift ratios are identical to the ones in Fig. 3 in the main text.
@@
Since we already know that our results reduce back to those in population genetics when $\alpha\ll 1$ ($\eta\approx 0)$, we now focus on the case of large $\eta$ and $\alpha\gg 1$. In this regime, due to the absolute value in $R$, the solutions to $R=1$ have different types of behavior depending on the value of $s$. Recall that we require $0\leq f\leq 1$. When $s_{\mathrm{inv}}<\sqrt{4\eta/N_{\mathrm{eff}}}$ (or $\tilde{s}_{\mathrm{inv}}<2/\alpha$), there is only one solution with $s_{\mathrm{inv}}-\eta f<0$: (Fig. S5a)
$f_{\mathrm{barrier}}\sim\frac{2}{N_{\mathrm{eff}}|s_{\mathrm{inv}}|}\left(\sqrt{1+\frac{4\eta}{N_{\mathrm{eff}}s_{\mathrm{inv}}^{2}}}-\mathrm{sgn}(s_{\mathrm{inv}})\right)^{-1}\,.$ (141)
When $s_{\mathrm{inv}}>\eta$, there is also only one solution but with $s_{\mathrm{inv}}-\eta f>0$: (Fig. S5c)
$f_{\mathrm{barrier}}\sim\frac{2}{N_{\mathrm{eff}}s_{\mathrm{inv}}}\left(\sqrt{1-\frac{4\eta}{N_{\mathrm{eff}}s_{\mathrm{inv}}^{2}}}+1\right)^{-1}\,.$ (142)
@@
Therefore, the dynamics of $f$ is qualitatively the same as the one in classic population genetics for $s_{\mathrm{inv}}<\sqrt{4\eta/N_{\mathrm{eff}}}$ or $s_{\mathrm{inv}}>\eta$, with $f_{\mathrm{barrier}}$ being the drift threshold.
@@
In contrast, the dynamics of $f$ is qualitatively distinct for $\sqrt{4\eta/N_{\mathrm{eff}}}<s_{\mathrm{inv}}\lesssim\eta$. We see that there are three solutions to $R=1$: (Fig. S5b)
$f_{\mathrm{barrier}}$ $=\frac{2}{N_{\mathrm{eff}}s_{\mathrm{inv}}}\left(1+\sqrt{1-\frac{4\eta}{N_{\mathrm{eff}}s_{\mathrm{inv}}^{2}}}\right)^{-1},$ (143)
$f_{\mathrm{coexist}}^{\mathrm{lower}}$ $=\frac{2}{N_{\mathrm{eff}}s_{\mathrm{inv}}}\left(1-\sqrt{1-\frac{4\eta}{N_{\mathrm{eff}}s_{\mathrm{inv}}^{2}}}\right)^{-1},$ (144)
$f_{\mathrm{coexist}}^{\mathrm{upper}}$ $=\frac{2}{N_{\mathrm{eff}}s_{\mathrm{inv}}}\left(\sqrt{1+\frac{4\eta}{N_{\mathrm{eff}}s_{\mathrm{inv}}^{2}}}-1\right)^{-1}\,$, (145)
with $f_{\mathrm{barrier}}<f_{\mathrm{coexist}}^{\mathrm{lower}}<f_{\mathrm{coexist}}^{\mathrm{upper}}$. Note that we have $f_{\mathrm{coexist}}^{\mathrm{lower}}\simeq f_{\mathrm{coexist}}^{\mathrm{upper}}\simeq f^{*}\simeq\tilde{s}_{\mathrm{inv}}$, hence these two frequencies correspond to the bounds of the coexistence region centered at $f^{*}$.
@@
As demonstrated in the main text, our results in fixation probability and mean absorption/extinction time can be qualitatively understood using the nonlinear behavior of $R$ (Fig. S5b). Suppose a mutant starts with $f=f_{0}\ll 1$ at the beginning. It must overcome the first barrier at $f=f_{\mathrm{barrier}}$ or it becomes extinct. After that, selection brings the mutant to the region $f_{\mathrm{coexist}}^{\mathrm{lower}}<f<f_{\mathrm{coexist}}^{\mathrm{upper}}$, which is the coexistence region. Note that the deterministic fixed point $f^{*}\simeq\tilde{s}_{\mathrm{inv}}$ is within this region. Since $\left<df/dt\right>>0$ when $f<f_{\mathrm{coexist}}^{\mathrm{lower}}$ and $\left<df/dt\right><0$ when $f>f_{\mathrm{coexist}}^{\mathrm{upper}}$, both boundaries of the coexistence region are reflecting and the mutant is trapped inside this region. It can escape this region and reach near the absorbing boundaries $f=0,1$ only when the demographic noise reaches the tail of its distribution with exponentially low probability.
@@
Now we see that if $s_{\mathrm{inv}}<s_{c}=\eta/2$, the coexistence region is closer to $f=0$ and it is exponentially more likely for the demographic noise to drive the mutant to extinction. Therefore, $p_{\mathrm{fix}}$ is exponentially suppressed for $s_{\mathrm{inv}}<s_{c}$. In contrast, when $s_{\mathrm{inv}}>s_{c}$, the coexistence region is closer to $f=1$ and it is exponentially more likely for the demographic noise to drive the mutant to fixation. The mutant becomes extinct mostly because it cannot overcome the first barrier at $f=f_{\mathrm{barrier}}$, hence $p_{\mathrm{fix}}$ agrees with Kimura’s formula for $s_{\mathrm{inv}}>s_{c}$.
@@
The above argument also explains the behavior of mean absorption/extinction time. It grows and decays exponentially in the region $\sqrt{4\eta/N_{\mathrm{eff}}}<s_{\mathrm{inv}}\lesssim\eta$ due to the required time to escape the coexistence region. More precisely, the time grows exponentially as $s_{\mathrm{inv}}$ increases and the coexistence region becomes farther from $f=0$. The growth continues till $s_{\mathrm{inv}}=s_{c}$. For $s_{\mathrm{inv}}>s_{c}$, the coexistence region becomes closer to fixation, and the mean absorption time decays exponentially as fixation becomes more likely. For the mean extinction time, although the time for escaping from the coexistence region to $f=0$ is even longer, it also becomes exponentially unlikely that the mutant becomes extinct after reaching the coexistence region. Namely, a mutant becomes extinct mainly due to the first barrier at $f=f_{\mathrm{barrier}}$. As a result, the mean extinction time reaches maximum at $s_{\mathrm{inv}}=s_{c}$ and decays exponentially for $s_{\mathrm{inv}}>s_{c}$.
Appendix H Mean extinction time
@@
For the regimes without the coexistence region, i.e. $s_{\mathrm{inv}}<\sqrt{4\eta/N_{\mathrm{eff}}}$ or $s_{\mathrm{inv}}>\eta$, the extinction time can be approximated by the time required to reach the barrier at $f=f_{R=1}$; see Eqs. (141) and (142). For $s_{\mathrm{inv}}<\sqrt{4\eta/N_{\mathrm{eff}}}$, the barrier is reflecting and the mutant becomes extinct in the same order of magnitude of time after reaching the barrier. For $s_{\mathrm{inv}}>\eta$, the barrier is absorbing and the mutant can no longer become extinct after passing through the barrier. Now, the time for reaching the barrier can be estimated using Eq. (134); since $\Delta t$ increases with $f$, the required time to reach $f=f_{R=1}$ is dominated by $\Delta t(f_{R=1})$. Therefore, for $s_{\mathrm{inv}}<\sqrt{4\eta/N_{\mathrm{eff}}}$, the mean extinction time $T$ is
$T\sim\frac{2}{|s_{\mathrm{inv}}|}\left(\sqrt{1+\frac{4\eta}{N_{\mathrm{eff}}s_{\mathrm{inv}}^{2}}}-\mathrm{sgn}(s_{\mathrm{inv}})\right)^{-1}\,$, (146)
and for $s_{\mathrm{inv}}>\eta$,
$T\sim\frac{2}{s_{\mathrm{inv}}}\left(\sqrt{1-\frac{4\eta}{N_{\mathrm{eff}}s_{\mathrm{inv}}^{2}}}+1\right)^{-1}\,.$ (147)
Both cases match the expectation from population genetics that $T\sim 1/|s_{\mathrm{inv}}|$ asymptotically.
On the other hand, the mean extinction time deviates from the expectation from population genetics when the coexistence region is present, i.e. when $\sqrt{4\eta/N_{\mathrm{eff}}}<s_{\mathrm{inv}}<\eta$. In this regime, the mean extinction time is dominated by the exponentially long time to escape the coexistence region driven by the stochastic dynamics. The scale of the escape time can be estimated by translating the stochastic dynamics into 1D diffusion in the presence of a potential $V(f)$. Since $f$ relaxes to near equilibrium around the deterministic fixed point $f^{*}\simeq\tilde{s}_{\mathrm{inv}}$ in the long term (but before being absorbed at $f=0,1$), the probability distribution $p(f)$ can be obtained from the Fokker-Planck equation:
$\frac{\partial p}{\partial t}$ $=-\frac{\partial}{\partial f}\left(f(1-f)(s_{\mathrm{inv}}-\eta f)p\right)+\frac{\partial^{2}}{\partial^{2}f}\left(\frac{f(1-f)}{2N_{\mathrm{eff}}}p\right)$ (148)
@@
$T_{0}\sim\exp\left(\frac{V(0)-V(\tilde{s}_{\mathrm{inv}})}{k_{B}T_{\mathrm{eff}}}\right)=\exp\left(\alpha^{2}\tilde{s}_{\mathrm{inv}}^{2}\right)\,.$ (153)
@@
$p_{0}\sim\frac{p(0)}{p(0)+p(1)}\sim\begin{cases}1&s<s_{c}\\ \frac{p(0)}{p(1)}=\exp\left(\alpha^{2}(1-\tilde{s}_{\mathrm{inv}})^{2}-\alpha^{2}\tilde{s}_{\mathrm{inv}}^{2}\right)&s>s_{c}\,.\end{cases}$ (154)
@@
$T\sim p_{0}T_{0}\sim\begin{cases}\exp\left(\alpha^{2}\tilde{s}_{\mathrm{inv}}^{2}\right)&s_{\mathrm{inv}}<s_{c}\\ \exp\left(\alpha^{2}(1-\tilde{s}_{\mathrm{inv}})^{2}\right)&s_{\mathrm{inv}}>s_{c}\,.\end{cases}$ (155)
In particular, the mean extinction time reaches maximum when $s_{\mathrm{inv}}=s_{c}\simeq\eta/2$, i.e.
@@
Appendix I Methods, including details of simulations
@@
After that, we randomly choose one of the surviving species with equal probability as the parent. Since we only focus on a single mutation, the distribution for which parent is chosen does not affect the analysis. Each time, we introduce a mutant associated with the chosen parent, sampled as described in Sec. A and B. We then solve the full differential equations with demographic noise using a generalized Euler method as explained below. Due to the multiplicative nature of the noise, the simulation can become numerically unstable if we sample the noise directly. Instead, we make use of the result in [77]. Within a time step $dt$, we can approximate that the total growth rates of the species remain unchanged, since the growth rates involve averages over all the species or resource abundances. In other words, the species dynamics within time $dt$ becomes
@@
To obtain fixation probabilities, we sample $10$ mutants with the same chosen parent. For each mutant, we adjust the value of $r_{m}$ (for generalized Lotka-Volterra model) or $m_{m}$ (for consumer-resource models) so that the invasion fitness is at a given value of $s_{\mathrm{inv}}$. We then run the above simulation $10^{3}$ times with different instances of demographic noise and count the number of fixations. We take the average of the above results to find the fixation probability $p_{\mathrm{fix}}(s_{\mathrm{inv}})$.
@@
Appendix J Parameters used in the figures
Throughout all figures, the initial mutant frequency is $f_{0}=0.09$. We set $A_{pm}=A_{mp}=\rho$ for all simulations of generalized Lotka-Volterra models.
@@
$S=50,\mu=2.0,\sigma=0.4,\mu_{r}=6.0,\sigma_{r}=0.6,\gamma=1,\rho=0.99,s_{\mathrm{inv}}=0.2,N_{p}=5.0,D=0.2\,.$ (160)
@@
$\sigma=0.4,\rho=0.9,N_{p}=2.6$ (162)
We change the control parameter $\alpha$ by using $D=1.0$ in Fig. 2(a) and $D=0.04$ in Fig. 2(b). In Fig. 2(e), we use
$\sigma=0.4,s_{\mathrm{inv}}=0,N_{p}=4.8\,.$ (163)
In Fig. 2(f), the simulation parameters for the case of low packing are the same as those for Fig. 2(b). To control the extent of packing, we keep all the parameters unchanged except using $\sigma=0.7$ for the case with high packing.
@@
In Figs. S3 and S4, we simulate generalized Lotka-Volterra model with parameters
$S=50,\mu=2.0,\sigma=0.4,\mu_{r}=6.0,\sigma_{r}=0.6,\gamma=1,D=0.2\,.$ (171)
Fig. S3 uses $s_{\mathrm{inv}}=0$ and Fig. S4 uses $s_{\mathrm{inv}}=0,0.4,0.8$. To obtain the exponentially suppressed fixation probabilities, we repeat the simulation for each mutant $10^{4}$ times to count the number of fixations.
In Fig. S5, the selection-drift ratios are identical to the ones in Fig. 3.